11. Calculate [OH- ] ions in a 0.125 M solution of nitrous acid. Is this solution acidic or basic; how do you know? Make sure you include the ionization equation. (6 marks)



12. Calculate the [H3O+] in a 2.00 L solution of NaOH, a strong base, if it contains 0.800 g of solute. Is this solution acidic or basic; how do you know? Make sure you include the dissociation equation. (6 marks)


1
Expert's answer
2022-05-25T07:33:04-0400

11. Calculate [OH] ions in a 0.125 M solution of nitrous acid. Is this solution acidic or basic; how do you know? Make sure you include the ionization equation.

Solution (11):

Nitrous acid (HNO2) is a weak acid (Ka = 4.5×10−4 for HNO2).

The balanced equation for the ionization of HNO2 is:

HNO2(aq) + H2O(l) ⇌ H3O+(aq) + NO2(aq)

Summarize the initial conditions, the changes, and the equilibrium conditions in the following ICE table:



Substitute the equilibrium concentrations into the expression for the acid ionization constant Ka:



Solving for x we get:

x = 0.00728

[H3O+] = x = 0.00728 M


Calculate the pH as the negative of the base-10 logarithm of [H3O+]:

pH = −log[H3O+] = −log(0.00728) = 2.14

pH = 2.14, it is less than 7 so this solution is acidic


For any aqueous solution at 25C:

[H3O+][OH] = Kw = 1.0×10−14

Therefore,

[OH] = Kw / [H3O+] = (1.0×10−14) / (0.00728) = 1.37×10−12

[OH] = 1.37×10−12 M


Answer (11): [OH] = 1.37×10−12 M; this solution is acidic




12. Calculate the [H3O+] in a 2.00 L solution of NaOH, a strong base, if it contains 0.800 g of solute. Is this solution acidic or basic; how do you know? Make sure you include the dissociation equation.

Solution (12):

The molar mass of sodium hydroxide (NaOH) is 40 g/mol

Therefore,

Moles of NaOH = (0.800 g NaOH) × (1 mol NaOH / 40 g NaOH) = 0.020 mol NaOH


Molarity = Moles of solute / Liters of solution

Therefore,

Molarity of NaOH solution = (0.020 mol) / (2.00 L) = 0.01 mol/L = 0.01 M


Sodium hydroxide (NaOH) is a strong base because it almost completely ionized in water.

The balanced equation for the ionization of NaOH is:

NaOH(aq) → Na+(aq) + OH(aq)

And thus a 0.01 M solution of NaOH is stoichiometric in Na+ and OH, i.e. [Na+] = [OH] = 0.01 M

[OH] = 0.01 M


For any aqueous solution at 25C:

[H3O+][OH] = Kw = 1.0×10−14

Therefore,

[H3O+] = Kw / [OH] = (1.0×10−14) / (0.01) = 1.0×10−12

[H3O+] = 1.0×10−12 M


Calculate the pH as the negative of the base-10 logarithm of [H3O+]:

pH = −log[H3O+] = −log(1.0×10−12) = 12

pH = 12, it is more than 7 so this solution is basic


Answer (12): [H3O+] = 1.0×10−12 M; this solution is basic